ABPO operational status:
The station operator visited the station and replaced the two Q330 digitizers and replaced the STS-2 with an STS-2.5 seismometer.
The station operator visited the station and replaced the STS-1s with a Trillium360SV seismometer.
Data quality improved for the STS-1N on day 2021,036 and for the STS-1V on day 2021,043. IDA engineers are still investigating the cause for the offsets, suspect some faults with the E300 electronics.
Data quality from the STS-1V and STS-1N is poor, showing excessive vertical offsets, as of day 2021,033-03:45 and 034-21:37, respectively.
The station operator upgraded the radio telemetry link, and installed two new station computers and a Paroscientific microbarometer.
The station operator replaced the STS-2 on day 2012,199-10:45 with one shipped from UCSD. Remote calibration is now possible, and the data quality has improved.
All channels from the STS-2 are showing abrupt changes in amplitudes as of day 2012,127-12:00 for reasons yet to be determined.
Telemetry to the station was restored after the station operator replaced the Freewave radios.
Telemetry to the station remains down. In addition to problems with the Freewave radios, there are communication problems between the Q330-10 digitizer and the ISP that are yet to be determined. IDA engineers are working with the station operator to diagnose and resolve these problems.
Telemetry to the station is down due to problems with the Freewave radio. A new radio will be shipped to the station. Data is still being acquired by IOGA.
An IDA engieer visited the station and upgraded the power system and telemetry link.
The station is back on-line as of day 2010,272-12:21 after an IDA engineer visited the station and upgraded it to a Quanterra Q330HR system. Seismometers include an STS-1, STS-2, and an episensor strong motion detector.
The station is down as of day 2010,265-01:54 due to problems with the ISP. An IDA engineer is enroute to upgrade the station.
The station was down from day 2009,305-13:00 to 2009,307-13:18 due to power outages. The station is currently on-line, however, UPS batteries are low and need to be recharged.
The station was down from about day 2008,267 to 2008,269 due to problems with power to the station after a recent storm. The station came back on-line after the station operator restored power to the station.
The station was down from about day 2008,100-09:29 to 2008,101-07:56 due to a power problem and problems with the connection to the station. The station came back on-line when power was restored and the station operator replaced the ethernet switch and cables.
The station was down from about day 2007,254-22:29 to 2007,257-07:35 due to problems with the DAS. Data acquisition resumed after the station operator cycled power to the DAS.
The vertical channels from the STS-1 seismometer are nonseismic as of about day 2007,244-09:07.
The station is back online as of about day 193-12:56 after the station operator reconfigured the cabling to the DAS.
The station is down as of about day 189-18:55 due to communication problems between the DAS and the ISP that are yet to be determined.
The vertical channels from the STS-1 seismometer had background noise until about day 127 due to a faulty cable connected to the DAS.
New station ABPO installed in Madagascar with a MK8 data acquisition system. Sensors installed include STS-1 broadband and STS-2 short period seismometers, a strong motion sensor, and a Paroscientific microbarograph.